Al can paint the house in 5 1/3 hours. It takes Ben 8 hours to do the same job. How long would it take if they worked together?
I've tried:
(1 / 5 1/3) + (1/8) = (1/t) and then I get stuck.

Answer by lwsshak3(11628)

Al can paint the house in 5 1/3 hours. It takes Ben 8 hours to do the same job. How long would it take if they worked together?
I've tried:
(1 / 5 1/3) + (1/8) = (1/t) and then I get stuck.
(1 /(16/3) + (1/8) = (1/t)
3/16+2/16=1/t
5/16=1/t
t=16/5=3 1/5
How long would it take if they worked together? 3 1/5 hrs